OpenStudy (anonymous):

trying to solve for x. In a polygon the four angles are: (20-5x)+(2x+80)+(x2-4)+(x2+5). [where x2 is x squared].

OpenStudy (anonymous):

polygons with four sides have angles that add to 360 degrees (20 - 5x) + (2x + 80) + (x^2 - 4) + (x^2 + 5) = 360 Add like terms 2x^2 - 3x + 101 = 360 Set the one side = to 0 2x^2 - 3x - 259 = 0 Do you know how to solve this?

OpenStudy (anonymous):

Are you sure your problem is accurate... It doesn't come out to be a "nice" number and I would think it would at this level.

OpenStudy (anonymous):

Yes, I checked the angles are accurate. I am not clear about how you got 2x^2 - 3x - 259 = 0. I thought it would be 2x^2+2x-264=0. Any suggestions?